# Searching for strong lensing by late-type galaxies in UNIONS

### v1

@@ 3.1 The challenges of finding rare objects
where $P(\text{L}\,|\,\text{C})$ is the probability that something classified as a lens candidate is indeed a lens, this represents the ratio of lenses over machine-selected candidates, and is often called the ‘purity’ of the lens search. On the other hand, $P(\text{C}\,|\,\text{L})$ is the probability that the CNN correctly classifies the image of a lens as a lens. This is also known as the true positive rate (TPR), or recall. $P(\text{L})$ is the probability of a random stamp being a lens, which is the prevalence of lensing on the sky. $P(\text{C})$ is the probability that the CNN classifies any given stamp as a lens. Since every stamp shows either a lens or not, we marginalise the last term as
@@
[…]
@@
[…]
@@ 3.2 The visual inspection
To estimate the number of galaxy-scale lenses in the parent sample, we conduct a systematic visual inspection of randomly selected sources. This procedure allows us to establish a baseline against which we can compare the performance of our CNN classifier. Additionally, it provides some estimates of the potential of the parent sample for future, more general, lens finding efforts. For this, we split the parent sample into four $r$-band bins of magnitude $(17,18]$, $(18,19]$, $(19,20]$, and $(20,20.5]$. We then randomly select 30 0003000030\,00030 000 sources per bin to be inspected, for a total of 120 000120000120\,000120 000 sources probed. A similar procedure is used in Acevedo Barroso et al. (2024) to estimate the number of lens candidates in the entire Euclid Wide Survey.
For the visual inspection, we follow a three-step methodology, in which the sources selected in one step go on to be reinspected in the next one:
@@ 1. Step 1:
    the experts individually inspect the sources in mosaics;
@@ 2. Step 2:
    the experts individually inspect the sources one at a time;
@@ 3. Step 3:
For this purpose, we have reimplemented the visualisation tools used in Rojas et al. (2022) and Savary et al. (2022) using the Qt6 framework. [^1] Available at https://github.com/ClarkGuilty/Qt-stamp-visualizer. The tools correspond to a mosaic viewer that shows multiple sources in a rectangular grid, and a 1-by-1 sequential tool that displays a single source at a time. Both tools allow the user to change the colour-map and scaling function used to display the stamps. The mosaic tool is used for binary classification: the user goes through pages of mosaics and clicks on the stamps that show hints of lensing. By contrast, the 1-by-1 sequential tool allows for the detailed classification of sources into the following non-overlapping categories:
@@
  • A, shows clear lensing features and no additional information is needed;
@@
  • B, shows lensing features but additional information is required to confirm its lensing nature;
@@
  • C, shows lensing features, but these can be explained by other phenomena;
  • X, definitively not a lens.
@@
Additionally, the 1-by-1 sequential tool offers quick access to three $grz$ colour-composite images from the DESI Legacy Imaging Survey (Dey et al. 2019, Legacy Survey henceforth): a colour-composite image covering the area of the stamp inspected ($12.256″12.256″($$\,\times\,$$)$; a colour-composite image displaying the environment around the stamp in a larger field of view ($2.13′2.13′($$\,\times\,$$)$; and a colour-composite covering the same area as the stamp, but displaying the residuals of the inference modelling used to construct the DESI source catalogue. We present an example of the two tools in Fig. 1, for which the 1-by-1 sequential tool is showing the environment around the inspected stamp.
@@
Figure 1: Visualisation tools used for the visual inspections. Left panel: the mosaic tool showing an $8\times 8$ grid, whilst highlighting sources classified as ‘lenses’ with a capitalised ‘L’ instead of the UNIONS stamp. Right panel: The 1-by-1 sequential tool showing a spiral galaxy on the left side, and the larger field of view from Legacy Survey on the right side. The source is classified as a ‘non-lens’ and further sub-classified as a ‘spiral’. Both applications are using logarithmic scale and a grey colour-map for the UNIONS stamps.
@@
The first two steps of the visual inspection were performed by J.A.A.B., B.C., and F.C. Each one inspected a different set of 10 0001000010\,00010 000 sources from each of the magnitude bins during the mosaic step. Then, all the candidates selected by any of the three human classifiers were pooled together for the individual 1-by-1 inspection. Lastly, the final group inspection was done by six expert classifiers J.A.A.B., B.C., F.C., C.L., R.G., and K.R. During the final inspection, the classifiers had access to $r$- and $u$-band UNIONS imaging, as well as the Legacy Survey data previously described. They also used, when available, $grizy$ imaging data from Pan-STARRS (Chambers et al. 2016; Flewelling et al. 2020), and spectral data from the the Sloan Digital Sky Survey (SDSS; Blanton et al. 2017).
Figure 2: Lens candidates discovered during the prevalence study, sorted by visual inspection grade and right ascension. We show a colour composite of the $r$- and $u$-bands when $u$-band data are available, otherwise we just display a grey-scale $r$-band image. The number in the bottom left is the CNN score, and the letter in the bottom right corresponds to the final human classification. We highlight the edge-on lenses with their name in cyan, and the candidates previously reported in SLED with an overlaid orange square.
@@ 3.3 Extrapolating to the whole parent sample
[…]
@@
A+B 17 to 18 328 345328345328\,345328 345 1 11 [8, 35] 5 55 [40, 90]
A+B 18 to 19 999 184999184999\,184999 184 0 < $<\,37$ 3 100 [70, 195]
A+B 19 to 20 2 848 36728483672\,848\,3672 848 367 0 < $<\,108$ 4 380 [271, 677]
A+B 20 to 20.5 2 802 38128023812\,802\,3812 802 381 0 < $<\,106$ 2 187 [129, 431]
C 17 to 18 328 345328345328\,345328 345 5 55 [40, 90] 11 120 [95, 166]
C 18 to 19 999 184999184999\,184999 184 3 100 [70, 195] 6 200 [148, 317]
C 19 to 20 2 848 36728483672\,848\,3672 848 367 1 95 [68, 311] 2 190 [131, 438]
C 20 to 20.5 2 802 38128023812\,802\,3812 802 381 0 < $<\,106$ 4 374 [267, 666]
We use Bayes’ theorem to estimate the number of lens candidates in the parent sample from the number of candidates found in the visual inspection. To do so, we make the following assumptions:
  1. 1.
    we have identified all the lenses in the sample probed (30 0003000030\,00030 000 per magnitude bin);
  2. 2.
    being randomly selected, the sample probed is representative of the entire parent sample.
Under these assumptions, the likelihood $P(k\,|\,K)$ of finding $k$ lens candidates in the probed sample out of a parent sample containing $K$ lens candidates is given by
@@
where $n$ is the number of sources probed and $N$ is the number of sources in the parent sample. Then, we estimate $P(K\,|\,k)$: the posterior probability that there are $K$ lenses in the parent sample, given that we have found $k$ in the sample probed. Following Bayes’ theorem: $P(K\,|\,k)\propto P(k\,|\,K)P(K)$. Last, we institute a prior on the number of lenses in the parent sample $P(K)$, such that the probability of a random source showing lensing features ($K/N$) is distributed uniformly logarithmically between $10^{-10}$ and unity.
We make the calculation independently for every magnitude bin, but join the grades A and B to increase the signal. The results are summarised in Table 1. We further classify 10 of the candidates as edge-on based on the morphology of their deflector galaxy. However, only one of them is deemed grade A, while the rest are deemed grade C. This good candidate belongs in the brightest magnitude bin, hinting at a core problem of searching for edge-on lenses: the lensed light is very often blended with the deflector, leading to uncertain classifications. The low signal in our measurement means that the estimates are prior-dominated for all the other magnitude bins and correspond only to upper bounds. Overall, we estimate at least eight grade A and B edge-on lens candidates, and at least 500 candidates for lensing in general. This is consistent with a prevalence rate of 1 lens for every 10 0001000010\,00010 000 sources for lensing in general, and 1 in 30 0003000030\,00030 000 for edge-on lenses.
@@ 4 Simulating a training data set
We use the data-driven methodology introduced by Cañameras et al. (2020) and reproduced by Rojas et al. (2022) and Savary et al. (2022) to produce a large data set of simulated lenses to train our CNN. This requires us to select a set of real UNIONS sources as potential deflector galaxies. For each deflector, we assume a composite mass model comprising both a baryonic component and a dark matter halo. We then simulate the lensing of real galaxies from deep HST imaging data, and add the resulting arcs on the UNIONS images of our selection of deflectors. We use the same HST sources with the Subaru Hyper Suprime-Cam (HSC) colours described in Rojas et al. (2022). For all the lensing calculations and image simulations we use the Python package Lenstronomy (Birrer & Amara 2018; Birrer et al. 2021), and we use Astropy (Astropy Collaboration et al. 2022) to handle the imaging data. Figure 3 gives a visual summary of the complete simulation algorithm. We describe below the different elements of the process.
Figure 3: Schematic description of our lens simulation procedure. To produce each simulated stamp, we start with a potential deflector and a target Einstein radius, $\theta_{\mathrm{E}}$. Then, we pick a set of parameters for a mass model consistent with the target $\theta_{\mathrm{E}}$ and the deflector’s light, and perform a check where we make sure that $\log M$, $f_{\ast}$, and $\Upsilon$ are within physical ranges. After that, we pick a source using the target $z_{s}$ provided by the mass model parameters, and place it randomly around the caustics before simulating its lensing by the mass model. Next, we convolve the lensed source image with the deflector’s PSF, then downsample to match the pixel size of CFIS, and add Poissonian noise. At this point, we add the original stamp of the deflector to get the final mock. Lastly, we perform a contrast check to make sure that the lensing features are visible. If it passes the check, we accept the mock and proceed to the next deflector. Otherwise, we reposition the source relative to the caustics and try a second time.
@@ 4.1 Selection of potential CFIS deflectors
We begin the process by selecting the edge-on late-type galaxies in our CFIS data. To do so, we apply an ellipticity cut similar to the one used by Sygnet et al. (2010), namely $0.920.6<e<$ 0.92, with
@@
[…]
@@ 4.2 Light and mass models
To simulate the mass distribution of deflector galaxies, we use a composite model where baryons and dark matter are treated independently. This approach is preferable to power-law or isothermal models because, for this type of galaxy, the bulge and disc components dominate the lensing mass over the rounder dark halo. Furthermore, the relative proportion between the bulge, disc, and dark halo has observable effects on the lensing features.
For the baryonic part of the composite model, we use the light profile as seen in the actual data, normalised by a constant stellar mass-to-light ratio, $\Upsilon$. For the dark halo component, we use a spherical Navarro-Frenk-White profile (NFW, see Navarro et al. 1996). Since the light profiles of the deflectors are highly elliptical, we approximate an elliptical Srsic profile as the difference between two cored elliptical isothermal profiles. This is known as a ‘chameleon’ profile, previously used in Dutton et al. (2011). In this work, we follow the parametrisation introduced by Suyu et al. (2014):
@@
where ${\xi}\equiv(\xi_{1},\xi_{2})=\@vec{\theta}-\@vec{\theta_{0}}$ is the angular position relative to the centroid of the profile, $q$ is the axis-ratio, and $L_{0}$ is the surface brightness at 11″1  ″ from the centroid along the major axis. In addition, $w_{\rm c}$ and $w_{\rm t}$ are the sizes of the two cored isothermal profiles that are subtracted from each other, with the additional condition that $w_{\rm t}>w_{\rm c}$.
Because in practice our deflector galaxies can display a prominent bulge component, we also use a ‘double chameleon’ profile to model their light distribution. The double chameleon is parametrised as the sum of two concentric chameleon profiles, each one with its own ellipticity and core sizes:
$L_{\rm d}\left(\@vec{\xi},\@vec{X_{1}},\@vec{X_{2}},r,L_{0}\right)=L_{\rm c}\left(\@vec{\xi},\@vec{X_{1}},\frac{L_{0}}{1+r}\right)+L_{\rm c}\left(\@vec{\xi},\@vec{X_{2}},\frac{L_{0}}{1+1/r}\right)\,,$ (8)
where the vector ${X_{\rm i}}\equiv(w_{\rm ic},w_{\rm it},q_{\rm i})$ corresponds to the parameters for the single chameleon $i$, and $r$ is the ratio between the profiles along the major axis at 11″1  ″ away from the common centroid. We use the chameleon profiles as proxies for the baryonic mass, by replacing the normalisation $L_{0}$ with $\alpha_{1}$, where $\alpha_{1}$ is now the deflection angle along the major axis at $1\arcsec$ away from the centroid. By contrast, the spherical NFW profile is parametrised in terms of its virial mass $M_{200}$ and concentration $c$. We use the scaling relations described in Duffy et al. (2008) to estimate the concentration $c$ for every $M_{200}$.
We fit the chameleon and double chameleon models to the light of every potential deflector using the r-band stamps, their corresponding PSF images, and a circular pixel mask with radius $4.5″$ around the centroid of the light distributions. To do so, we use a log-likelihood that combines the standard pixel-wise noise-normalised square error with penalties for ellipticities larger than $e>0.98$ and for negative fluxes. We minimise the log-likelihood using first a particle swarm optimiser (Kennedy & Eberhart 1995) with 200 particles and 50 iterations, and then downhill simplex optimisation (Nelder & Mead 1965) with a maximum of 3000 iterations. Then, we select between the chameleon and double chameleon models using the reduced $\chi^{2}$. Unsurprisingly, deflectors with prominent bulges are better modelled by the double chameleon, while for bulgeless discs the single chameleon suffices. Good models easily reach a reduced $\chi^{2}$ close 1, since we have reliable estimates of the pixel-wise noise via the RMS maps provided by the CFIS reduction pipeline. Finally, we discard deflectors with reduced $\chi^{2}>3$. The whole process leaves us with a total of 558 199558199558\,199558 199 potential deflectors with good light models.

### v2

@@ 3.1 The challenges of finding rare objects
where $P(\text{L}\,|\,\text{C})$ is the probability that something classified as a lens candidate is indeed a lens, this represents the ratio of lenses over machine-selected candidates, and is often called the ‘purity’ of the lens search. On the other hand, $P(\text{C}\,|\,\text{L})$ is the probability that the CNN correctly classifies the image of a lens as a lens. This is also known as the true positive rate (TPR), or recall. $P(\text{L})$ is the probability of a random stamp being a lens, which is the prevalence of lensing on the sky. $P(\text{C})$ is the probability that the CNN classifies any given stamp as a lens. Since every stamp shows either a lens or not, we marginalised the last term as
@@
with $P(\text{C}\,|\,\text{Non-L})$ being the probability that the network mistakenly classifies an image that does not contain any lensing features as a lens. This is known as the false positive rate (FPR). Equation 1 becomes
@@
Taking the conservative estimate that one object shows lensing features for every 10 00010\,000 galaxies, then $P(\text{L})=10^{-4}$. If we further assume a network capable of correctly classifying all lenses in a sample and 99% of the non-lenses, then $P(\text{C}\,|\,\text{L})=1$ and $P(\text{C}\,|\,\text{Non-L})=0.01$. From there, Eq. 3 implies that only $1\%$ of the stamps classified as lenses by the network are in fact real lenses, that is, $P(\text{L}\,|\,\text{C})\approx 0.01$.
@@ 3.2 The visual inspection
To estimate the number of galaxy-scale lenses in the parent sample, we conduct a systematic visual inspection of randomly selected sources. This procedure allows us to establish a baseline against which we can compare the performance of our CNN classifier. Additionally, it provides some estimates of the potential of the parent sample for future, more general, lens finding efforts. For this, we split the parent sample into four $r$-band bins of magnitude $(17,18]$, $(18,19]$, $(19,20]$, and $(20,20.5]$. We then randomly select 30 00030\,000 sources per bin to be inspected, for a total of 120 000120\,000 sources probed. A similar procedure is used in Acevedo Barroso et al. (2025) to estimate the number of lens candidates in the entire Euclid Wide Survey.
For the visual inspection, we followed a three-step methodology in which the sources selected in one step go on to be reinspected in the next one:
@@ 1. Step 1:
    The experts individually inspect the sources in mosaics,
@@ 2. Step 2:
    The experts individually inspect the sources one at a time, and
@@ 3. Step 3:
For this purpose, we reimplemented the visualisation tools used in Rojas et al. (2022) and Savary et al. (2022) using the Qt6 framework. [^1] Available at https://github.com/ClarkGuilty/Qt-stamp-visualizer. The tools correspond to a mosaic viewer that shows multiple sources in a rectangular grid, and a one-by-one sequential tool that displays a single source at a time. Both tools allow the user to change the colour-map and scaling function used to display the stamps. The mosaic tool is used for binary classification: the user goes through pages of mosaics and clicks on the stamps that show hints of lensing. By contrast, the one-by-one sequential tool allows for detailed classification of sources into the following non-overlapping categories:
@@
  • A: The source shows clear lensing features and no additional information is needed.
@@
  • B: The source shows lensing features but additional information is required to confirm its lensing nature.
@@
  • C: The source shows lensing features, but these can be explained by other phenomena.
  • X: The source is definitively not a lens.
Grade A candidates clearly display multiple images or lensed arcs following typical lensing configurations. By contrast, grade B candidates may exhibit potential multiple images but lack typical lensing symmetries, usually spectroscopy or high-resolution imaging is enough to confirm these candidates lenses. Grade C candidates still resemble simulated mock lenses, but the imaging can also be explained by random alignment of galaxies or a merger event. Because of this, grade C candidates often require both high-resolution imaging and spectroscopy for confirmation. Variations of this classification scheme are routinely used in lens finding works (e.g., Rojas et al. 2022; Savary et al. 2022; Acevedo Barroso et al. 2025; Nagam et al. 2025; Euclid Collaboration: Walmsley et al. 2025; Euclid Collaboration: Lines et al. 2025).
@@
Additionally, the one-by-one sequential tool offers quick access to three grzgrz colour-composite images from the DESI Legacy Imaging Survey (Dey et al. 2019, Legacy Survey henceforth): a colour-composite image covering the area of the stamp inspected (12.256×12.256)($\times\,$$$); a colour-composite image displaying the environment around the stamp in a larger field of view (2.13×2.13)($\times\,$$$); and a colour-composite covering the same area as the stamp, but displaying the residuals of the inference modelling used to construct the DESI source catalogue. We present an example of the two tools in Fig. 1, for which the one-by-one sequential tool is showing the environment around the inspected stamp.
@@
Figure 1: Visualisation tools used for the visual inspections. Left panel: Mosaic tool showing an $8\times 8$ grid whilst highlighting sources classified as ‘lenses’ with a capitalised ‘L’ instead of the UNIONS stamp. Right panel: One-by-one sequential tool showing a spiral galaxy on the left side and the larger field of view from Legacy Survey on the right side. The source is classified as a ‘non-lens’ and further sub-classified as a ‘spiral’. Both applications are using logarithmic scale and a grey colour-map for the UNIONS stamps.
@@
The first two steps of the visual inspection were performed by J.A.A.B., B.C., and F.C. Each inspector examined a different set of 10 00010\,000 sources from each magnitude bin during the mosaic step. Since none of the known edge-on lens candidates have UNIONS/CFIS observations, we used simulated mock lenses as a references (see Sects. 4 and D). All the candidates selected by any inspector were pooled together for individual one-by-one inspection. Finally, all candidates that received at least one grade A or grade B vote were reinspected in the final group inspection, which involved six expert classifiers: J.A.A.B., B.C., F.C., C.L., R.G., and K.R. After inspection and discussion of each candidate, the final grade was decided by a majority vote (i.e. at least four out of six votes). For this stage, the inspectors had access to $r$- and $u$-band UNIONS imaging, as well as the previously described Legacy Survey data. When available, they also used grizygrizy imaging data from Pan-STARRS (Chambers et al. 2016; Flewelling et al. 2020), and spectroscopy from the Sloan Digital Sky Survey (SDSS; Blanton et al. 2017).
@@ 3.3 Extrapolating to the whole parent sample
Table 1: Estimate of the number of lens candidates in the parent sample.  [^2] Information derived from the expert visual inspection of n=30 000n=\$30\,000\$ randomly selected sources per magnitude bin. Uncertainties are given for a 68% confidence interval. Results are shown for both strong lensing in general, and for edge-on lenses only. If no candidates were found, the estimation corresponds to a prior-dominated upper bound. The names of the columns in parenthesis match the elements of Eq. 4.
@@
A+B 17 to 18 328 345328\,345 1 11 [8, 35] 5 55 [40, 90]
A+B 18 to 19 999 184999\,184 0 < 37<\,37 3 100 [70, 195]
A+B 19 to 20 2 848 3672\,848\,367 0 < 108<\,108 4 380 [271, 677]
A+B 20 to 20.5 2 802 3812\,802\,381 0 < 106<\,106 2 187 [129, 431]
C 17 to 18 328 345328\,345 5 55 [40, 90] 11 120 [95, 166]
C 18 to 19 999 184999\,184 3 100 [70, 195] 6 200 [148, 317]
C 19 to 20 2 848 3672\,848\,367 1 95 [68, 311] 2 190 [131, 438]
C 20 to 20.5 2 802 3812\,802\,381 0 < 106<\,106 4 374 [267, 666]
We use Bayes’ theorem to estimate the number of lens candidates in the parent sample from the number of candidates found in the visual inspection. To do so, we made the following assumptions: (1) We have identified all the lenses in the sample probed (30 00030\,000 per magnitude bin), and (2) being randomly selected, the sample probed is representative of the entire parent sample. With these assumptions, the likelihood P(k|K)P(k\,|\,K) of finding $k$ lens candidates in the probed sample out of a parent sample containing $K$ lens candidates is given by
@@
where $n$ is the number of sources probed and $N$ is the number of sources in the parent sample. Then, we estimate P(K|k)P(K\,|\,k): the posterior probability that there are $K$ lenses in the parent sample, given that we have found $k$ in the sample probed. Following Bayes’ theorem: $P(K\,|\,k)\propto P(k\,|\,K)P(K)$. Last, we institute a prior on the number of lenses in the parent sample $P(K)$, such that the probability of a random source showing lensing features ($K/N$) is distributed uniformly logarithmically between $10^{-10}$ and unity.
Figure 2: Lens candidates discovered during the prevalence study sorted by visual inspection grade and right ascension. We show a colour composite of the $r$- and $u$-bands when the $u$-band data are available, otherwise we just display a grey scale $r$-band image. The number in the bottom left is the CNN score, and the letter in the bottom right corresponds to the final human classification. We highlight the edge-on lenses with their name in cyan, and the candidates previously reported in SLED with an overlaid orange square.
We made the calculation independently for every magnitude bin, but join the grades A and B to increase the signal. The results are summarised in Table 1. We further classify 10 of the candidates as edge-on based on the morphology of their deflector galaxy. However, only one of them is deemed grade A, while the rest are deemed grade C. This good candidate belongs in the brightest magnitude bin, hinting at a core problem of searching for edge-on lenses: the lensed light is very often blended with the deflector, leading to uncertain classifications. The low signal in our measurement means that the estimates are prior-dominated for all the other magnitude bins and correspond only to upper bounds. Overall, we estimate at least eight grade A and B edge-on lens candidates, and at least 500 candidates for lensing in general. This is consistent with a prevalence rate of 1 lens for every 10 00010\,000 sources for lensing in general, and 1 in 30 00030\,000 for edge-on lenses.
@@ 4 Simulating a training dataset
We used the data-driven methodology introduced by Cañameras et al. (2020) and reproduced by Rojas et al. (2022) and Savary et al. (2022) to produce a large dataset of simulated lenses to train our CNN. This required us to select a set of real UNIONS sources as potential deflector galaxies. For each deflector, we assumed a composite mass model comprising both a baryonic component and a dark matter halo. We then simulated the lensing of real galaxies from deep HST imaging data and added the resulting arcs on the UNIONS images of our selection of deflectors. We used the same HST sources with the Subaru Hyper Suprime-Cam (HSC) colours described in Rojas et al. (2022). For all the lensing calculations and image simulations, we used the Python package Lenstronomy (Birrer and Amara 2018; Birrer et al. 2021), and we used Astropy (Astropy Collaboration et al. 2022) to handle the imaging data. Figure 3 gives a visual summary of the complete simulation algorithm. We describe the different elements of the process in the following sub-sections.
Figure 3: Schematic description of our lens simulation procedure. To produce each simulated stamp, we started with a potential deflector and a target Einstein radius, $\theta_{\mathrm{E}}$. Then, we picked a set of parameters for a mass model consistent with the target $\theta_{\mathrm{E}}$ and the deflector’s light, and we performed a check where we made sure that $\log M$, $f_{\ast}$, and $\Upsilon$ are within physical ranges. After that, we picked a source using the target $z_{s}$ provided by the mass model parameters, and we placed it randomly around the caustics before simulating its lensing by the mass model. Next, we convolved the lensed source image with the deflector’s PSF, downsampled to match the pixel size of CFIS, and added Poissonian noise. At this point, we added the original stamp of the deflector to get the final mock. Lastly, we performed a contrast check to make sure that the lensing features are visible. If it passed the check, we accepted the mock and proceeded to the next deflector. Otherwise, we repositioned the source relative to the caustics and tried a second time.
@@ 4.1 Selection of potential CFIS deflectors
We began the process by selecting the edge-on late-type galaxies in our CFIS data. To do so, we applied an ellipticity cut similar to the one used by Sygnet et al. (2010), namely $0.6<e<0.92$, with
@@
Then, we cross-match the selected galaxies with the Sloan Digital Sky Survey (SDSS) data release 17 (Abdurro’uf et al. 2022) to obtain photometric redshifts (photo-$z$), K-corrections, and distance moduli. We keep only those galaxies with $\rm{\texttt{photoErrorClass}}\,=\,1$ in the SDSS catalogue, ensuring the best possible estimation of the photo-$z$ error. We note that the SDSS data do not cover the entire CFIS footprint; however, this is not a problem as long as there are enough galaxies with photo-$z$ to build a training dataset. On the other hand, the photo-$z$ error increases with magnitude, such that it is about six times larger for the faintest galaxies ($\Delta z\approx 0.13$) than for the brightest ones ($\Delta z\approx 0.02$). Since the photo-$z$ is used to estimate the distance to the deflectors, and from there the luminosity, total mass, and mass-to-light ratio of the mock lenses, a large error in the redshift estimation can introduce significant uncertainties on the physical properties of the mocks. Fortunately, the relative error in photo-$z$ for almost all potential deflectors is between $10\%$ to $30\%$ which is low enough to be acceptable for the generation of mocks. Thus, only the catastrophic photo-$z$ estimations introduce significant errors in the physical parameters of the mock lenses.
Finally, we removed bad deflectors, which we defined as galaxies that, based on their r-band image, would require a mass-to-light ratio larger than 12 to generate a lens with Einstein radius of $1\theta_{\mathrm{E}}$=\$\$, assuming a source at redshift $z=3.0$ and a single isothermal ellipsoid (SIE) model for the mass. We took $1\theta_{\mathrm{E}}$=\$\$ because it is almost twice the average seeing, guaranteeing the visibility of the lensing features. In total, we had 649 340649\,340 galaxies useful as potential deflectors. For consistency with the K-corrections and photo-$z$s, we used the same cosmology as SDSS for all relevant calculations, adopting flat $\Lambda\mathrm{CDM}$ with $\Omega_{\rm m}=0.2739$ and $H_{\rm 0}=$70.5\text{\,}\mathrm{km}\text{\,}{\mathrm{s}}^{-1}\text{\,}{\mathrm{Mpc}}^{-1}$\$.
@@ 4.2 Light and mass models
To simulate the mass distribution of deflector galaxies, we used a composite model where baryons and dark matter are treated independently. This approach is preferable to power-law or isothermal models because, for this type of galaxy, the bulge and disc components dominate the lensing mass over the rounder dark halo. Furthermore, the relative proportion between the bulge, disc, and dark halo has observable effects on the lensing features.
For the baryonic part of the composite model, we used the light profile as seen in the actual data, normalised by a constant stellar mass-to-light ratio, $\Upsilon$. For the dark halo component, we used a spherical Navarro-Frenk-White profile (NFW, see Navarro et al. 1996). Since the light profiles of the deflectors are highly elliptical, we approximated an elliptical Sérsic profile as the difference between two cored elliptical isothermal profiles. This is known as a ‘chameleon’ profile, previously used in Dutton et al. (2011). In this work, we followed the parametrisation introduced by Suyu et al. (2014):
@@
where $\@vec{\xi}\equiv(\xi_{1},\xi_{2})=\@vec{\theta}-\@vec{\theta_{0}}$ is the angular position relative to the centroid of the profile, $q$ is the axis-ratio, and $L_{0}$ is the surface brightness at 1 from the centroid along the major axis. In addition, $w_{\rm c}$ and $w_{\rm t}$ are the sizes of the two cored isothermal profiles that are subtracted from each other, with the additional condition that $w_{\rm t}>w_{\rm c}$.
Because in practice our deflector galaxies can display a prominent bulge component, we also used a ‘double chameleon’ profile to model their light distribution. The double chameleon is parametrised as the sum of two concentric chameleon profiles, each one with its own ellipticity and core sizes:
[…]
where the vector $\@vec{X_{\rm i}}\equiv(w_{\rm ic},w_{\rm it},q_{\rm i})$ corresponds to the parameters for the single chameleon $i$, and $r$ is the ratio between the profiles along the major axis at 1 away from the common centroid. We used the chameleon profiles as proxies for the baryonic mass, by replacing the normalisation $L_{0}$ with $\alpha_{1}$, where $\alpha_{1}$ is now the deflection angle along the major axis at $1\arcsec$ away from the centroid. By contrast, the spherical NFW profile is parametrised in terms of its virial mass $M_{200}$ and concentration $c$. We used the scaling relations described in Duffy et al. (2008) to estimate the concentration $c$ for every $M_{200}$.
We fitted the chameleon and double chameleon models to the light of every potential deflector using the r-band stamps, their corresponding PSF images, and a circular pixel mask with radius 4.5 around the centroid of the light distributions. For this, we used a log-likelihood that combined the standard pixel-wise noise-normalised square error with penalties for ellipticities larger than $e>0.98$ and for negative fluxes. We minimised the log-likelihood using first a particle swarm optimiser (Kennedy and Eberhart 1995) with 200 particles and 50 iterations, and then downhill simplex optimisation (Nelder and Mead 1965) with a maximum of 3000 iterations. Then, we selected between the chameleon and double chameleon models using the reduced $\chi^{2}$. Unsurprisingly, deflectors with prominent bulges were better modelled by the double chameleon, while for bulgeless discs the single chameleon sufficed. Good models easily reached a reduced $\chi^{2}$ close 1, since we have reliable estimates of the pixel-wise noise via the RMS maps provided by the CFIS reduction pipeline. Finally, we discarded deflectors with reduced $\chi^{2}>3$. The whole process left us with a total of 558 199558\,199 potential deflectors with good light models.